OpenStudy (anonymous):

Please Help With Algebra??? WILL GIVE A MEDAL AND WILL FAN!!! In a classroom, 5 out of 13 students surveyed said they had two or more siblings. If the entire school has 520 students, predict the number of students that have two or more siblings. Write a proportion to model the problem. Use x for your variable.

OpenStudy (anonymous):

Is it a written answer or is it a multiple choice

OpenStudy (anonymous):

it is a written answer question @mayekeroom

OpenStudy (aripotta):

ok so you have to use a proportion\[\frac{ part }{ whole }\frac{ 5 }{ 13 }=\frac{ x }{ 520 }\]

OpenStudy (aripotta):

cross-multiply and solve for x

OpenStudy (anonymous):

i don't really know how to do that?

OpenStudy (aripotta):

5(520) = 13x

OpenStudy (aripotta):

see what i did?

OpenStudy (anonymous):

after I solve 5 times 520 I divide both sides by 13?

OpenStudy (aripotta):

yes

OpenStudy (anonymous):

it's 200

OpenStudy (aripotta):

yep

OpenStudy (anonymous):

so the written proportion is \[\frac{ 5 }{ 13 }=\frac{ x }{ 520}\]

OpenStudy (aripotta):

yes

OpenStudy (anonymous):

Thank you @AriPotta :)

OpenStudy (aripotta):

\(\Large \bf \color{#008AFF}{n}\color{#0073FF}{o} ~\color{#005CFF}{p} \color{#0045FF}{r} \color{#0017FF}{o} \color{#0200FF}{b} \color{#3000FF}{l} \color{#4700FF}{e} \color{#7500FF}{m} \color{#8C00FF}{o}~\LARGE\color{#D100FF}{\ddot\smile}\)